In 1950, the Hong Kong Government built a hospital for about $500,000. In 1990, the county restored the courthouse for about $1.5 million. A price index for nonresidential construction was 20 in 1950, 80 in 1990, and 120 in 2015. According to these numbers, the hospital cost about _____ in 2015 dollars, which is _____ than the cost of the courthouse restoration in 2015 dollars.
Answer #1

Cost of the hospital in 2015 dollars = 500000*120/20 = 3000000 = 3 million

Cost of the courthouse restoration in 2015 = 1500000*120/80 = 2250000 = 2.25 million

so, the cost of the hospital is more than the cost of the restoration of the courthouse